OpenStudy (anonymous):

6ab^2/5a^3b divided by 12ab/15a^5b I dont know if im supposed to flip the second half and make the division into a multiplication or simplify first.

OpenStudy (anonymous):

Alright, first let's cut down the question into 2 parts :):\[\frac{6ab^2}{5a^3b}\] this part is divide by : \[\frac{12ab}{15a^5b}\] now let's just write it down in a simpler form ^_^: \[\frac{6ab^2}{5a^3b} \div \frac{12ab}{15a^5b}\] now there's a rule that says, when you divide 2 fractions or any other number, you have to flip the other fraction/number and multiply instead of divide like this: \[\frac{6ab^2}{5a^3b} \times \frac{15a^5b}{12ab}\] now let's simplify :) \[= \frac{3a^2b}{2}\] is it clearer now :)?

OpenStudy (anonymous):

thanks, so does that mean the flipping rule always comes first no matter what?

OpenStudy (anonymous):

yep ^_^ and np